OpenStudy (anonymous):

1 + 2/x = 3/x^2 Can someone solve this using factorising and quadratic formula pliz?

OpenStudy (mertsj):

\[1+\frac{2}{x}=\frac{3}{x^2}\] \[x^2+2x=3\] \[x^2+2x-3=0\]

OpenStudy (mertsj):

Can you factor that?

OpenStudy (anonymous):

ok sweet fanx mate for help

OpenStudy (mertsj):

yw

OpenStudy (anonymous):

@timkainavaka he multiplied throughout by the least common multiple of all of their denominators, \(x^2\).

OpenStudy (anonymous):

fanx oldrin